# Efficient Generative Model Training via Embedded Representation Warmup

### v2

@@ 5 Experiments
Table 1: System-level comparison on ImageNet 256$\times$256 with CFG. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. Results marked with an asterisk (*) use advanced CFG scheduling techniques; specifically, for our method, we apply the guidance interval scheduling from (Kynkäänniemi et al., 2024).
@@ 5.1 Setup
$\bullet$ Implementation Details. We adhere closely to the experimental setups described in DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024), unless otherwise noted. Specifically, we utilize the ImageNet dataset (Deng et al., 2009), preprocessing each image to a resolution of $\times 256$ 256 pixels. Following the protocols of ADM (Dhariwal & Nichol, 2021), each image is encoded into a compressed latent vector $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE (Rombach et al., 2022). For our model configurations, we employ the B/2 and XL/2 architectures as introduced in the SiT papers, which process inputs with a patch size of 2. To ensure a fair comparison with SiT models and REPA, we maintain a consistent batch size of 256 throughout training. Further experimental details, including hyperparameter settings and computational resources, are provided in Appendix C.
$\bullet$ Evaluation. We report Fréchet inception distance (FID; Heusel et al. 2017), sFID (Nash et al., 2021), inception score (IS; Salimans et al. 2016), precision (Pre.) and recall (Rec.) (Kynkäänniemi et al., 2019) using 50K samples. We also include CKNNA (Huh et al., 2024) as discussed in ablation studies. Detailed setups for evaluation metrics are provided in Appendix D.
@@
Table 2: FID comparisons with SiT-XL/2. In this table, we report the FID of ERW with SiT-XL/2 on ImageNet $\times 256$ 256 at various Training iterations. This table shows results from the full training phase, starting from a checkpoint pre-trained with our warmup phase. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar iterations or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without Classifier-Free Guidance.
@@
$\bullet$ Baselines. We use several recent diffusion-based generation methods as baselines, each employing different inputs and network architectures. Specifically, we consider the following four types of approaches: (a) Pixel diffusion: ADM (Dhariwal & Nichol, 2021), VDM+++++ + (Kingma & Gao, 2024), Simple diffusion (Hoogeboom et al., 2023), CDM (Ho et al., 2022), (b) Latent diffusion with U-Net: LDM (Rombach et al., 2022), (c) Latent diffusion with transformer+U-Net hybrid models: U-ViT-H/2 (Bao et al., 2023), DiffiT (Hatamizadeh et al., 2024), and MDTv2-XL/2 (Gao et al., 2023), and (d) Latent diffusion with transformers: MaskDiT (Zheng et al., 2024), SD-DiT (Zhu et al., 2024), DiT (Peebles & Xie, 2023), and SiT (Ma et al., 2024). Here, we refer to Transformer+U-Net hybrid models that contain skip connections, which are not originally used in pure transformer architecture. Detailed descriptions of each baseline method are provided in   Appendix E.
5.2 Comparison to Prior Methods
Table 1 summarizes our results on ImageNet $\times 256$ 256 under Classifier-Free Guidance (CFG). Our ERW significantly boosts the convergence speed of SiT-XL/2, enabling strong FID scores at 350 epochs.   Figure 3 illustrates generated samples on ImageNet, further confirming the high-quality outputs achieved by ERW.
5.3 Effect of ERW: Quantitative Gains and Training Efficiency
Performance without Classifier-Free Guidance.
We begin by examining how ERW influences SiT-XL/2’s FID when no CFG is used.
  1. (a)
    Efficient FID Improvements. In  Table 5.1, ERW consistently achieves competitive or superior FID values compared to baselines. For instance, ERW reaches an FID of 12.1 with 100k warmup + 100k full training, markedly outperforming the REPA method (Yu et al., 2024) within the same iteration budget.
  2. (b)
    Leveraging Pretrained Features. This gain highlights the advantage of injecting pretrained semantic priors via warmup, thereby accelerating convergence in the subsequent full diffusion training.
Summary. By incorporating a short warmup step, ERW enables the model to focus its capacity on generative decoding rather than learning semantic features from scratch, resulting in improved FID at fewer training steps.
Warmup versus full training.
Next, we analyze how splitting the total training budget between warmup and full diffusion training impacts both generation quality and computational overhead. As shown in  Figure 4, FLOPs for warmup is fewer than the full training.
Table 3: Analysis of ERW on ImageNet 256$\times$256. All models are SiT-B/2 trained for 50K iterations. All metrics except FID without Classifier-Free Guidance. We fix $\lambda=0.5$ here. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively.
[…]
MoCov3-B 8 NT-Xent 61.1 7.6 22.38 0.42 0.58
MoCov3-L 8 NT-Xent 73.0 8.0 17.96 0.38 0.52
CLIP-L 8 NT-Xent 58.9 7.7 23.68 0.44 0.54
DINOv2-B 8 NT-Xent 55.6 7.8 25.45 0.44 0.56
DINOv2-L 8 NT-Xent 55.5 7.8 25.45 0.44 0.56
DINOv2-g 8 NT-Xent 59.4 7.6 25.53 0.44 0.56
Summary. A dedicated, relatively brief warmup phase materializes substantial downstream speedups and quality gains, confirming that investing in warmup alignment is a cost-effective strategy.
5.4 Ablation Studies
Target representation.
Table 4: Analysis of ERW places influences in SiT-B/2. All models are based on SiT-B/2 and trained for 50K iterations under the batch size of 256 without using Classifier-Free Guidance on ImageNet $\times 256$ 256. $\downarrow$ indicates lower values are better. Results empirically validate our hypothesis that placing ERW at the forefront of the architecture yields optimal performance.
[…]
SiT-B/2 + REPA  (Yu et al., 2024) 78.2 11.71 17.1 0.33 0.48
DINOv2-B 0-8 NT-Xent 54.2 8.12 27.2 0.45 0.59
DINOv2-B 1-9 NT-Xent 69.1 13.0 18.7 0.37 0.51
DINOv2-B 2-10 NT-Xent 67.7 13.4 19.0 0.38 0.52
DINOv2-B 3-11 NT-Xent 67.5 11.8 19.5 0.38 0.52
DINOv2-B 4-11 NT-Xent 67.8 13.1 19.0 0.38 0.52
  1. (a)
    Universality of Pretrained Encoders. All encoders tested offer improvements over baselines, indicating that ERW can broadly benefit from a variety of strong representation models.
  2. (b)
    Marginal Differences among DINOv2 Variants. DINOv2-B, DINOv2-L, and DINOv2-g yield comparable gains, suggesting that ERW does not require the largest possible teacher encoder for effective representation transfer.
@@ 5.4 Ablation Studies
Table 5: Analysis of ERW depth, projection depth, and different dynamic or consistent projection loss $\lambda$ influences in SiT-XL/2. All models are based on SiT-XL/2 and trained for 100K iterations under a batch size of 256 without using Classifier-Free Guidance on ImageNet $\times 256$ 256. The target representation model is DINOv2-B, and the objective is NT-Xent. $\downarrow$ indicates lower values are better. The results show that a projection depth of 14 and a projection loss $\lambda$ of 4.0 yield substantial improvements in both FID and sFID, indicating an optimal configuration for model performance.
@@
Summary. Regardless of which high-quality self-supervised model is used, early-stage alignment provides a consistent boost in sample fidelity and training efficiency.
@@
  1. (a)
    Empirical Validation. In  Table 5.4, initializing the earliest layers (0–8) notably outperforms re-initializing middle or late sections (FID 54.2 vs. >67>67> 67).
  2. (b)
    Consistent with Circuit Perspective. This corroborates our three-stage diffusion circuit ( Section 3), underscoring that aligning deeper layers for representation can be suboptimal since those layers focus more on generative refinement.
Summary. Targeting the front portion of the network for warmup is crucial, reinforcing our theoretical claim that representation learning predominantly resides in early layers.
@@
We also investigate how the final projection head depth and the alignment-loss coefficient $\lambda$ affect training ( Table 5.4). The projection head, $\mathcal{T}_{\mathbf{\bm{\theta}}}$, is a deep MLP that, similar to REPA, maps intermediate transformer features to the dimensionality of the target representation.
  1. (a)
    Optimal Hyperparameters. Using 5 warmup layers, a projection head at depth 18, and $\lambda=0.5$ achieves an FID of 12.1 at 100k iterations—a substantial gain over baselines.
  2. (b)
    Trade-off in $\lambda$. Larger $\lambda$ offers stronger representation alignment initially but may disrupt convergence if pushed to extremes, highlighting the need for moderate scheduling.
Summary. Properly balancing projection depth and alignment strength is key to maximizing ERW’s benefits in practice.
@@
  1. (a)
    Initial Dip, Subsequent Recovery. Alignment falls early on as the pretrained features adjust to the diffusion objectives, but it then recovers and improves.
  2. (b)
    Role of Decaying Guidance. A decaying weight in the alignment term ( Section 4.2) fosters stable synergy between semantic alignment and generative refinement.
Summary. Representation alignment follows a U-shaped trajectory, revealing the model’s gradual reconfiguration of pretrained features for denoising tasks before distilling them into robust, generative-friendly embeddings.
@@
  1. (a)
    Improved Semantic Alignment. Compared to random initialization, ERW yields systematically higher CKNNA scores, confirming stronger semantic preservation.
  2. (b)
    Evolving Layer-wise Semantics. The alignment initially drops then recovers, mirroring the trends seen in  Figure 5 and supporting the conclusion that pretrained features are effectively integrated rather than merely overwritten.
Summary. CKNNA measurements reinforce our findings that ERW preserves and refines semantic information across layers, enabling more discriminative feature representations to emerge during diffusion training.
5.5 Hyperparameter tuning
  1. (a)
  2. (b)
    To evaluate performance quickly, we reduce the sampling steps from the usual 250250250 to 505050 and generate only $10\mathrm{k}$ samples (instead of $50\mathrm{k}$) to compute a preliminary FID score.
@@ 6 Conclusion and Future Work
In this work, we introduced Embedded Representation Warmup (ERW), a novel two-phase training framework that significantly enhances the training efficiency of diffusion models. By dedicating an initial phase to align the model’s early layers with a pretrained encoder, ERW establishes a strong semantic foundation that accelerates the subsequent generative training. Our key innovation is the explicit separation of representation alignment and generation, which, when combined with a decaying alignment schedule, proves more effective than continuous, single-phase regularization. We demonstrated empirically that ERW leads to substantial speedups in training convergence—up to 11.5$\times$ compared to REPA and achieves FID=1.41 with 350 epochs. Our ablations confirmed that targeting the early layers is crucial and that the two-phase approach is a cost-effective strategy for high-fidelity generative modeling.
Limitations and Future Work.
Despite its effectiveness, our work has limitations. First, the theoretical decomposition of the score function, while motivationally useful, is an idealization; the practical implementation relies on heuristic choices for the two-phase schedule and alignment targets. Second, the performance gains are sensitive to hyperparameters like the warmup duration and projection head depth, which currently require a manual tuning process. Future work will explore dynamically adjusting the alignment schedule based on training progress, rather than using a fixed decay.
Appendix A Loss Function Decomposition Analysis
This section presents a conceptual decomposition of the diffusion model’s loss function. Important note: This analysis is an idealization and relies on strong assumptions about representation sufficiency and orthogonality. In a real neural network, the components are not perfectly orthogonal, and the decomposition serves as a conceptual framework to motivate our method rather than a rigorous mathematical proof.
Let the diffusion model’s score matching objective be defined as:

### v3

@@ 4 Experiments
Table 1: System-level comparison on ImageNet $\times256$ with CFG. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. Results marked with an asterisk (*) use advanced CFG scheduling techniques; specifically, for our method, we apply the guidance interval scheduling from (Kynkäänniemi et al., 2024).
@@ 4.1 Setup
$\bullet$ Implementation Details. We adhere closely to the experimental setups described in DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024), unless otherwise noted. Specifically, we utilize the ImageNet dataset (Deng et al., 2009), preprocessing each image to a resolution of $256\times 256$ pixels. Following the protocols of ADM (Dhariwal & Nichol, 2021), each image is encoded into a compressed latent vector $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE (Rombach et al., 2022). For our model configurations, we employ the B/2 and XL/2 architectures as introduced in the SiT papers, which process inputs with a patch size of 2. To ensure a fair comparison with SiT models and REPA, we maintain a consistent batch size of 256 throughout training. Further experimental details, including hyperparameter settings and computational resources, are provided in Appendix D.
$\bullet$ Evaluation. We report Fréchet inception distance (FID; Heusel et al. 2017), sFID (Nash et al., 2021), inception score (IS; Salimans et al. 2016), precision (Pre.) and recall (Rec.) (Kynkäänniemi et al., 2019) using 50K samples. We also include CKNNA (Huh et al., 2024) as discussed in ablation studies. Detailed setups for evaluation metrics are provided in Appendix E.
@@
Table 2: FID comparisons with SiT-XL/2. In this table, we report the FID of ERW with SiT-XL/2 on ImageNet $256\times 256$ at various Training iterations. Here is only full training without warmup, because we load a well trained warmuped checkpoint. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar iterations or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without Classifier-Free Guidance.
@@
$\bullet$ Baselines. We use several recent diffusion-based generation methods as baselines, each employing different inputs and network architectures. Specifically, we consider the following four types of approaches: (a) Pixel diffusion: ADM (Dhariwal & Nichol, 2021), VDM++++ (Kingma & Gao, 2024), Simple diffusion (Hoogeboom et al., 2023), CDM (Ho et al., 2022), (b) Latent diffusion with U-Net: LDM (Rombach et al., 2022), (c) Latent diffusion with transformer+U-Net hybrid models: U-ViT-H/2 (Bao et al., 2023), DiffiT (Hatamizadeh et al., 2024), and MDTv2-XL/2 (Gao et al., 2023), and (d) Latent diffusion with transformers: MaskDiT (Zheng et al., 2024), SD-DiT (Zhu et al., 2024), DiT (Peebles & Xie, 2023), and SiT (Ma et al., 2024). Here, we refer to Transformer+U-Net hybrid models that contain skip connections, which are not originally used in pure transformer architecture. Details are provided in   Appendix F.
@@
Table 3: Analysis of ERW depth, projection depth, and different dynamic or consistent projection loss $\lambda$ influences in SiT-XL/2. All models are based on SiT-XL/2 and trained for 100K iterations under a batch size of 256 without using Classifier-Free Guidance on ImageNet $256\times 256$. The target representation model is DINOv2-B, and the objective is NT-Xent. $\downarrow$ indicates lower values are better. The results show that a projection depth of 14 and a projection loss $\lambda$ of 4.0 yield substantial improvements in both FID and sFID, indicating an optimal configuration for model performance.
@@
4.2 Comparison
Table 1 summarizes our results on ImageNet $256\times 256$ under Classifier-Free Guidance (CFG). Our ERW significantly boosts the convergence speed of SiT-XL/2, enabling strong FID scores at just 350 epochs. As shown in Table 1, our method achieves an FID of 1.41 in 350 epochs, that REPA requires 800 epochs to approach, demonstrating a high speedup while achieving state-of-the-art performance.   Figure 3 illustrates generated samples, further confirming the high-quality outputs achieved by ERW.
4.3 ERW Efficiency
We begin by how ERW influences SiT-XL/2’s FID when w/o CFG.
$\bullet$ Efficient FID Improvements. In  Table 4.1, ERW consistently achieves competitive or superior FID values compared to baselines. For instance, ERW reaches an FID of 12.1 with 100k warmup + 100k full training, markedly outperforming the REPA method (Yu et al., 2024) which scores 19.4 within the same budget.
$\bullet$ Leveraging Pretrained Features. This gain highlights the advantage of injecting pretrained semantic priors via warmup, thereby accelerating the full training.
Warmup versus full training.
Table 4: Analysis of ERW on ImageNet $\times256$. All models are SiT-B/2 trained for 50K iterations. All metrics except FID without Classifier-Free Guidance. We fix $\lambda=0.5$ here. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively.
[…]
MoCov3-B 8 NT-Xent 61.1 7.6 22.38 0.42 0.58
MoCov3-L 8 NT-Xent 73.0 8.0 17.96 0.38 0.52
CLIP-L 8 NT-Xent 58.9 7.7 23.68 0.44 0.54
DINOv2-B 8 NT-Xent 55.6 7.8 25.45 0.44 0.56
DINOv2-L 8 NT-Xent 55.5 7.8 25.45 0.44 0.56
DINOv2-g 8 NT-Xent 59.4 7.6 25.53 0.44 0.56
Next, we analyze how splitting the total training budget between warmup and full diffusion training impacts both generation quality and computational overhead. As shown in  Figure 4, the FLOPs for the warmup phase are significantly lower than for the full training phase.
4.4 Ablation Studies
Table 5: Analysis of ERW places influences in SiT-B/2. All models are based on SiT-B/2 and trained for 50K iterations under the batch size of 256 without using Classifier-Free Guidance on ImageNet $256\times 256$. $\downarrow$ indicates lower values are better. Results empirically validate our hypothesis that placing ERW at the forefront of the architecture yields optimal performance.
[…]
SiT-B/2 + REPA  (Yu et al., 2024) 78.2 11.71 17.1 0.33 0.48
DINOv2-B 0-8 NT-Xent 54.2 8.12 27.2 0.45 0.59
DINOv2-B 1-9 NT-Xent 69.1 13.0 18.7 0.37 0.51
DINOv2-B 2-10 NT-Xent 67.7 13.4 19.0 0.38 0.52
DINOv2-B 3-11 NT-Xent 67.5 11.8 19.5 0.38 0.52
DINOv2-B 4-11 NT-Xent 67.8 13.1 19.0 0.38 0.52
Target representation.
$\bullet$ Universality of Pretrained Encoders. All encoders tested offer improvements over baselines, indicating that ERW can benefit from a range of representation models.
$\bullet$ Marginal Differences among DINOv2 Variants. DINOv2-B, DINOv2-L, and DINOv2-g yield comparable gains, suggesting that ERW does not require the largest possible teacher encoder for effective representation transfer. This suggests that ERW is not limited to a specific encoder architecture but can leverage a wide range of powerful, pretrained feature extractors, making it a versatile tool for accelerating diffusion model training.
@@ 4.4 Ablation Studies
$\bullet$ Empirical Validation. In  Table 4.4, initializing the earliest layers (0–8) notably outperforms re-initializing middle or late sections (FID 54.2 vs. >67>67).
$\bullet$ Consistent with Circuit Perspective. This corroborates our three-stage diffusion circuit ( Section 3), underscoring that aligning deeper layers for representation can be suboptimal since those layers focus on generation. Targeting the initial layers for warmup is therefore crucial, reinforcing our theoretical claim that representation learning is predominantly the function of the early network stages, while later stages are specialized for generative refinement.
@@
We also investigate how the final projection head depth and the alignment-loss coefficient $\lambda$ affect training ( Table 4.1). The projection head, $\mathcal{T}_{\mathbf{\bm{\theta}}}$, is a deep MLP that maps the features to the dimensionality of the target representation as same as REPA.
$\bullet$ Empirical Validation. Using 5 warmup layers, a projection head at depth 18, and $\lambda=0.5$ achieves an FID of 12.1 at 100k iterations—a substantial gain over baselines.
$\bullet$ Trade-off in $\lambda$. Larger $\lambda$ offers stronger representation alignment initially but may disrupt convergence if pushed to extremes, highlighting the need for moderate scheduling.
@@
$\bullet$ Initial Dip, Subsequent Recovery. Alignment falls early on as the pretrained features adjust to the diffusion objectives, but it then recovers and improves.
$\bullet$ Role of Decaying Guidance. A decaying weight in the alignment term ( Section 3.3) fosters stable synergy between semantic alignment and generative refinement. The representation alignment thus follows a U-shaped trajectory, revealing the model’s initial adaptation of pretrained features to the diffusion task, followed by a distillation into robust, generation-aligned embeddings.
@@
$\bullet$ Improved Semantic Alignment. ERW yields systematically higher CKNNA scores, confirming stronger semantic preservation.
$\bullet$ Evolving Layer-wise Semantics. The alignment initially drops then recovers, mirroring the trends seen in  Figure 5 and pretrained features are effectively integrated rather than merely overwritten.
@@ 5 Conclusion and Future Work
[…]
Appendix A LLM Usage Statement
LLMs were used solely as auxiliary tools for grammar checking and language polishing. They did not contribute to the generation of research ideas, the design of experiments, the development of methodologies, data analysis, or any substantive aspects of the research. All scientific content, conceptual contributions, and experimental results are entirely the work of the authors. The authors take full responsibility for the contents of this paper.
Appendix B Theoretical Analysis
In this section, we provide a principled theoretical foundation for our ERW. We move beyond the empirical intuition of entangled objectives and demonstrate that our approach naturally emerges from a more fundamental perspective: conditional score matching within an augmented probability space. This formulation recasts the generative modeling problem as one where semantic understanding is an explicit conditional variable, thereby justifying the decoupling of representation learning from the synthesis process.
B.1 Preliminaries
To ensure clarity, we first establish the key mathematical objects used in our analysis. Let $\mathcal{X}$ be the high-dimensional data space (e.g., images), $\mathcal{Z}$ be the compressed latent space from a VAE, and $\mathcal{R}$ be the semantic representation space. We work with the following variables:
  • $\mathbf{x}\in\mathcal{X}$: A sample from the data distribution $p_{\text{data}}(\mathbf{x})$.
  • $\mathbf{z}_{0}\in\mathcal{Z}$: The clean latent representation of $\mathbf{x}$, obtained via a VAE encoder $\mathcal{H}_{\bm{\theta}_{\text{VAE}}}(\mathbf{x})$.
  • $\mathbf{z}_{t}\in\mathcal{Z}$: The noisy latent at time $t\in[0,1]$, defined by the forward process $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
  • $\mathbf{r}\in\mathcal{R}$: A high-level semantic representation vector corresponding to $\mathbf{x}$.
The key functions and models in our framework are:
  • $\mathbf{f}_{\text{rep}}:\mathcal{X}\to\mathcal{R}$: A powerful, pretrained, and fixed representation model (e.g., DINOv2) that maps a clean image $\mathbf{x}$ to its semantic representation $\mathbf{r}$.
  • $\mathbf{F}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{r})$: The diffusion model we aim to train, which predicts the score or velocity, potentially conditioned on a semantic representation $\mathbf{r}$.
  • $\mathcal{R}_{\bm{\theta}_{\text{L2R}}}$: The sub-network corresponding to the L2R circuit, which extracts representations from $\mathbf{z}_{t}$.
  • $\mathcal{G}_{\bm{\theta}_{\text{R2G}}}$: The sub-network corresponding to the R2G circuit, which performs generation based on an extracted representation.
The goal of a diffusion model is to learn the score function $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})$, which guides the reverse process from noise back to data. In the standard formulation, this requires learning to denoise across all time steps $t\in[0,1]$ without explicit semantic guidance.
B.2 A Principled View via an Augmented Probability Space
We formalize the intuition of decoupling representation and generation by constructing an augmented probability space that explicitly includes the semantic representation $\mathbf{r}$ as a random variable. This principled view demonstrates that our two-phase training strategy naturally emerges from optimizing a conditional score-matching objective in this richer probabilistic landscape.
B.2.1 Construction of the Augmented Space
We define an augmented probability space over the tuple $(\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{r},t)$, where the joint distribution factorizes as:
$p(\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{r},t)=p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,p(\mathbf{r}\mid\mathbf{z}_{0})\,p(\mathbf{z}_{0})\,p(t)$ (15)
This factorization leverages the conditional independence assumptions inherent in the diffusion process. Specifically, given the clean latent $\mathbf{z}_{0}$, the noisy latent $\mathbf{z}_{t}$ is independent of the semantic representation $\mathbf{r}$, and both are independent of the time variable $t$. Each component has a clear interpretation:
$p(\mathbf{z}_{0})$ $=\int p_{\text{data}}(\mathbf{x})\delta(\mathbf{z}_{0}-\mathcal{H}_{\bm{\theta}_{\text{VAE}}}(\mathbf{x}))d\mathbf{x}$ (16)
$p(t)$ $=\mathcal{U}[0,1]\quad\text{(uniform time distribution)}$ (17)
$p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)$ $=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ (18)
$p(\mathbf{r}\mid\mathbf{z}_{0})$ $=\delta\left(\mathbf{r}-\mathbf{f}_{\text{rep}}\left(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0})\right)\right)$ (19)
Here, $\mathcal{D}_{\bm{\theta}_{\text{VAE}}}$ denotes the VAE decoder that maps latents back to image space. The distribution $p(\mathbf{z}_{0})$ in  (16) represents the VAE’s learned prior over clean latents, induced by the data distribution through the encoder. The forward kernel (18) follows the standard diffusion forward process with noise scheduling parameters $\alpha_{t}$ and $\sigma_{t}$.
The key insight is that equation (19) deterministically links semantic representations to clean latents through the Dirac delta function, transforming the unconditional generation problem into a semantically-conditioned one. This constraint ensures that every clean latent $\mathbf{z}_{0}$ has a uniquely associated semantic representation $\mathbf{r}$, creating a deterministic mapping from the latent space to the representation space.
B.2.2 Marginal and Conditional Distributions
From the joint distribution, we can derive several important marginal and conditional distributions through careful integration.
Marginal over noisy latents: The marginal distribution over noisy latents is obtained by integrating out the semantic representation $\mathbf{r}$:
$p(\mathbf{z}_{t},t)$ $=\int\int\int p(\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{r},t)\,d\mathbf{z}_{0}\,d\mathbf{r}$ (20)
$=\int\int\int p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,p(\mathbf{r}\mid\mathbf{z}_{0})\,p(\mathbf{z}_{0})\,p(t)\,d\mathbf{z}_{0}\,d\mathbf{r}$ (21)
$=p(t)\int p(\mathbf{z}_{0})p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\left(\int p(\mathbf{r}\mid\mathbf{z}_{0})d\mathbf{r}\right)d\mathbf{z}_{0}$ (22)
$=p(t)\int p(\mathbf{z}_{0})p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,d\mathbf{z}_{0}$ (23)
where the integral $\int p(\mathbf{r}\mid\mathbf{z}_{0})d\mathbf{r}=1$ since $p(\mathbf{r}\mid\mathbf{z}_{0})$ is a valid probability distribution. This recovers the standard marginal distribution used in unconditional diffusion models.
Joint marginal over $(\mathbf{z}_{t},\mathbf{r},t)$: More critically for our analysis, we can compute the joint marginal over $(\mathbf{z}_{t},\mathbf{r},t)$ by integrating out only $\mathbf{z}_{0}$:
$p(\mathbf{z}_{t},\mathbf{r},t)$ $=\int p(\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{r},t)\,d\mathbf{z}_{0}$ (24)
$=\int p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,p(\mathbf{r}\mid\mathbf{z}_{0})\,p(\mathbf{z}_{0})\,p(t)\,d\mathbf{z}_{0}$ (25)
$=p(t)\int p(\mathbf{z}_{0})p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\delta\left(\mathbf{r}-\mathbf{f}_{\text{rep}}\left(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0})\right)\right)d\mathbf{z}_{0}$ (26)
Using the property of the Dirac delta function, this integral evaluates to:
$p(\mathbf{z}_{t},\mathbf{r},t)$ $=p(t)\int_{\mathbf{z}_{0}:\mathbf{f}_{\text{rep}}(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0}))=\mathbf{r}}p(\mathbf{z}_{0})p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,d\mathbf{z}_{0}$ (27)
where the integration is restricted to the set of clean latents $\mathbf{z}_{0}$ that produce the semantic representation $\mathbf{r}$ when decoded and passed through the representation function.
Conditional distribution for generation: We can also derive the conditional distribution of clean latents given noisy latents and semantic representations:
$p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)$ $=\frac{p(\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{r},t)}{p(\mathbf{z}_{t},\mathbf{r},t)}$ (28)
$=\frac{p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,p(\mathbf{r}\mid\mathbf{z}_{0})\,p(\mathbf{z}_{0})\,p(t)}{p(\mathbf{z}_{t},\mathbf{r},t)}$ (29)